A force of 5N gives a mass m1 an acceleration of 10m/s2 and mass m2 an acceleration of 20m/s
What acceleration would it give if both masses are tied together?

Grade:9

1 Answers

Arun
6 years ago
Dear Yash
 
When a force of 5N gives an acceleration of 10ms-² to a body 
then its mass = 5/10 kg = 0.5kg
so, m₁ = 0.5kg

when a force of 5N gives an acceleration of 20ms⁻² to a body 
then its mass = 5/20 kg = 0.25kg
so, m₂ =  0.25kg
when the two bodies r tied total mass of the system bcomes = 0.75kg
so acceleration of the system will b = force/mass = 5/0.75 = 6.67 ms⁻²
 
Regards
Arun (askIITians forum expert)